# Robotic Policy Learning via Human-assisted Action Preference Optimization

### v2

@@ Abstract
Establishing a reliable and iteratively refined robotic system is essential for deploying real-world applications. While Vision-Language-Action (VLA) models are widely recognized as the foundation model for such robotic deployment, their dependence on expert demonstrations hinders the crucial capabilities of correction and learning from failures. To mitigate this limitation, we introduce a Human-assisted Action Preference Optimization method named HAPO, designed to correct deployment failures and foster effective adaptation through preference alignment for VLA models. This method begins with a human-robot collaboration framework for reliable failure correction and interaction trajectory collection through human intervention. These human-intervention trajectories are further employed within the action preference optimization process, facilitating VLA models to mitigate failure action occurrences while enhancing corrective action adaptation. Specifically, we propose an adaptive reweighting algorithm to address the issues of irreversible interactions and token probability mismatch when introducing preference optimization into VLA models, facilitating model learning from binary desirability signals derived from interactions. Through combining these modules, our human-assisted action preference optimization method ensures reliable deployment and effective learning from failure for VLA models. The experiments conducted in simulation and real-world scenarios prove superior generalization and robustness of our framework across a variety of manipulation tasks.
[…]
@@ 1 Introduction
Ensuring safe interactions in unconstrained environments while fostering continuous improvement is crucial for the development of robust robotic manipulation systems in real-world scenarios [10, 26, 43]. Benefiting from the capacity for generalizable reasoning and scalable learning, Vision-Language-Action (VLA) models [3, 4, 8, 18, 41, 30] have been widely recognized as the foundation model for such robotic deployment systems. However, their performance in achieving high field-ready success rates in unconstrained, unpredictable real-world environments remains a significant limitation. This discrepancy presents a key challenge: how to integrate these developing Vision-Language-Action models into practical scenarios?
@@
To ensure reliable deployment and continuous learning ability of robotic systems, interactive imitation learning frameworks [7, 17] have been developed to refine error-prone trajectories via iterative human-in-the-loop correction feedback. Among them, behavior cloning [15, 21, 33] has been widely utilized to fine-tune a base policy model with manually corrected intervention data in a supervised learning fashion. In contrast, recent methods [20, 24] seek to propose effective off-policy reinforcement learning algorithms from sub-optimal human intervention trajectories. However, behavior cloning fails to fully exploit failure trajectories, which are valuable signals for learning robust policies. At the same time, reinforcement learning methods encounter significant scalability limitations in training large-scale VLA models, due to the inherent instability and challenge of developing generalizable value functions. To date, the effective adaptation of VLA models for downstream manipulation tasks remains understudied, particularly within sub-optimal human intervention paradigms.
To enable reliable deployment and stable learning from interactions, we propose a Human-assisted Action Preference Optimization method named HAPO, for autoregressive VLA models. This method integrates two critical components: the human-robot collaboration framework for reliable deployment, and the action preference optimization process for iterative improvement of VLA models. As shown in Figure 1(a), the human-robot collaboration deployment framework allows real-time human interventions during policy execution, ensuring reliable task completion when the robot encounters challenging situations. To mitigate the proportion imbalance of corrective action, we propose a balanced sampling method to provide proportional representation of interaction data for further VLA preference optimization. As shown in Figure 1(b), we introduce the action preference optimization process to fully leverage the sub-optimal interaction trajectories for stable VLA models optimization, which helps avoid failure actions and encourages the adoption of corrective actions. Through iterative human-robot collaboration deployment and action preference optimization process, our method can continuously enhance the VLA model’s capabilities via environment interaction, ensuring sustained improvements in performance and adaptability to dynamic downstream manipulation tasks.
@@
Figure 1: Pipeline of human-assisted action preference optimization method. This method consists of two components: (a) a human-robot collaboration deployment framework for reliable deployment and interaction trajectory collection with human intervention. (b) an action preference optimization process with adaptive reweighting for VLA models learning from sub-optimal human intervention trajectories. The size of circles represents their weight during training.
@@
However, there exist two challenges in adopting preference alignment optimization on autoregressive VLA models fine-tuning: (1) Irreversible interaction: While LLMs require paired preference data for optimization, the irreversible interaction process makes it difficult to gather meaningful paired positive-negative action samples under the same observation conditions. (2) Token probability mismatch: Autoregressive VLA models discretize continuous actions into tokens, causing a mismatch between token probabilities and true prediction loss, complicating preference alignment in robot action tokens prediction. To address these problems, we first employ Kahneman & Tversky’s prospect theory [11, 38] to formulate a preference alignment objective function that learns from binary desirability signals derived from interaction. This objective relaxes the demands of preference pairs, making it particularly suitable for learning from irreversible robotic interaction trajectories. Furthermore, we propose an adaptive reweighting method that leverages decoded continuous actions to guide preference optimization in the discrete action token space. This approach addresses the challenge of action token probability mismatch via the dynamic modulation of sample-wise training weighting, thereby concentrating gradient optimization on failure-prone interaction actions. Through weight refinement, we apply preference alignment optimization to VLA models, enhancing performance and stability when learning from sub-optimal manipulation correction trajectories.
@@
To systematically evaluate the effectiveness of our proposed system, we conduct comprehensive experiments in RoboMimic [27] simulation environments. The empirical results demonstrate that our HAPO method facilitates rapid adaptation in in-distribution scenarios while maintaining robust performance across a variety of unseen perturbations. Furthermore, lifelong learning experiments demonstrate the framework’s capacity for iterative improvement through human intervention. To evaluate the practical viability of the proposed framework, we conducted real-world experiments on fine-grained packing-hole tasks under a range of disruption conditions, demonstrating its robustness and applicability in real-world robotic manipulation scenarios.
2 Related Work
@@ 2.1 Vision-Language-Action Models
Achieving generalizable robotic manipulation remains a significant challenge within the field of robotics. Motivated by recent advances in foundation models [19, 37, 40, 42, 45], some works [6, 12, 39] attempt to construct large-scale real-world robotic datasets to facilitate the development of generalizable Vision-Language Action (VLA) models. Building upon these datasets, recent research [5, 18, 31] formulates robotic action prediction as a next token prediction problem within the framework of VLMs. In contrast, alternative studies [23, 36] investigate the applicability of diffusion-based methods to model multi-modal action distributions, thereby facilitating robustness in manipulation tasks. While these works focus on behavior cloning from expert demonstrations, Grape [46] proposes a trajectory-level preference alignment method to boost generalizability by incorporating both successful and failed trials. However, the requirements of paired trajectories under the same conditions make it infeasible in real-world scenarios. In this work, we build a human-assisted action preference optimization method to ensure reliable deployment in unconstrained robotic manipulation scenarios, alongside the preference alignment method that continuously refines the VLA model performance by systematically integrating human-in-the-loop intervention data.
@@ 2.2 Preference Alignment of Large Language Models
Contemporary methods [9, 28, 29, 47] implement Reinforcement Learning from Human Feedback (RLHF) through a two-stage method, which first trains a reward estimation model and optimizes LLMs to maximize the given estimated reward with a reinforcement learning method [13, 34]. However, this paradigm is slow and unstable in practice. DPO [32] proposes a single-step alternative that reparameterizes the RLHF objective into a closed-form loss function to directly maximize the log-likelihood margin between preferred and dispreferred outputs. Extending this framework, KTO [11] introduces the human-aware losses for learning from a binary signal of whether an output is desirable, which bypasses the need for intricate preference annotation altogether. In this work, we adapt the preference alignment optimization method for Vision-Language-Action models. Through an adaptive reweighting approach, we mitigate the irreversible interactions and token probability mismatch challenges when transferring preference learning methods from LLMs to VLA models.
@@ 2.3 Human-robot Interactive Learning
Interactive imitation learning [1, 7] has been proposed to refine robot actions through human feedback. While prior research [17, 21, 33] necessitates constant human supervision to intervene in the robot’s actions, more recent studies [14, 22, 44] have introduced dynamic models for automatic failure detection and real-time monitoring. In contrast, RLIF [24] leverages human intervention signals as rewards for off-policy RL, while HIL-SERL [25] presents a human-in-the-loop, vision-based RL system tailored for dexterous manipulation tasks. However, these RL approaches encounter difficulties in large-scale VLA model training, primarily due to unstable gradient optimization. In this work, we propose a human-assisted action preference optimization method to ensure the reliable deployment of robotics systems and the stable optimization of policies.
3 Human-assisted Action Preference Optimization
In this section, we introduce the human-assisted action preference optimization method as shown in algorithm 1 to enable reliable deployment and enhance continuous iterative improvement for the Vision-Language-Action (VLA) model. This framework comprises two components: a human-robot collaboration deployment framework to support real-time human intervention for reliable task execution, and an action preference alignment optimization process with adaptive reweighting for stable autoregressive VLA models fine-tuning through human interventions.
@@ 3.1 Human-robot Collaboration Deployment
3.2 Action Preference Optimization
Algorithm 1 Human-assisted Action Preference Optimization
@@ 3.2 Action Preference Optimization
2:$\mathcal{D}_{e}$: expert demonstrations, $\mathcal{D}_{h}$: interaction dataset, $\pi$: interaction policy
@@
6:$\theta^{*}\leftarrow\arg\max_{\theta}\mathbb{E}_{(o,a)\sim\mathcal{D}^{e}}\left[\log\pi^{0}_{\theta}(a|o)\right]$
7:$\mathcal{D}_{h}^{0}\xleftarrow{}\mathcal{D}_{e}$
@@
27:function Optimization($\pi_{\theta}$,$\pi_{ref}$,$\mathcal{D}_{h}$) 28:     for n gradient steps do 29:         Balanced Sample $(o_{i},a_{i},c_{i})_{i=1}^{i=B}\sim\mathcal{D}_{h}$ 30:         $l_{i}\xleftarrow{}|\pi_{\theta}(o_{i})-a_{i}|_{1}$ 31:         $w_{i}\xleftarrow{}\frac{l_{i}}{\sum_{i=1}^{i=B}l_{i}}$ 32:         if $c_{i}\neq 0$ then $\lambda_{D_{i}}=1-e^{-\beta_{D}*w_{i}}$ 33:         else $\lambda_{U_{i}}=e^{-\beta_{U}*w_{i}}$ 34:         end if 35:         $\theta^{*}=\arg\max_{\theta}\mathbb{E}[-v(o,a,\pi_{\theta},\pi_{ref},\lambda_{D},\lambda_{U})]$ 36:     end for 37:     return $\pi_{\theta^{*}}$ 38:end function
@@
In conclusion, we propose a human-assisted action preference optimization method as demonstrated in algorithm 1. This approach leverages the human-robot collaboration deployment for reliable task execution and interaction trajectories collection, while the action preference optimization process provides stable autoregressive VLA optimization with adaptive reweighting. Through iterative human-robot collaboration deployment and action preference optimization, we could achieve continual improvement from interaction with environments for autoregressive VLA models.
@@ 4 Experiments
To comprehensively evaluate our human-assisted action preference optimization method for effective downstream adaptation, we propose experiments to validate the following questions:
  • Could our method promote effective adaptation to in-distribution scenarios? Section 4.2
  • Could our method perform effectively in novel scenarios with various disruptions? Section 4.3
  • Could our method achieve iterative improvement during deployment? Section 4.4
  • Could our method generalize to various VLA models? Scetion 4.5
  • Could our method be applied in fine-grained real-world scenarios? Section 4.6
@@ 4.1 Experiment Settings
Implementation Details. In this work, we fine-tune the OpenVLA [39] model for target manipulation tasks as the base model. We employ LoRA [16] for parameter-efficient tuning, configuring rank $r=32$ with a batch size of 16 across 8 NVIDIA A100 GPUs. Further, we deploy the base model to interact with environments, where human operators perform real-time corrective interventions via a SpaceMouse device to rectify failures during execution. Based on the interaction trajectories collected during task execution, we fine-tune the base model $\pi_{ref}$ with our action preference optimization method, using a learning rate of $-55e{\text{-}}$ 5 and a batch size of 8 across 4 NVIDIA A100 GPUs. We set $K=10$ to automatically identify and annotate undesirable behaviors. To ensure the stability of preference alignment training, we employ balanced sampling to ensure that each batch contains 50% expert actions, 25% human intervention actions, and 25% failure actions.
Simulation Environments Details. We conduct experiments on RoboMimic [27] simulation. In the RoboMimic environment, we fine-tune the pretrained OpenVLA model for 4 long-horizon manipulation tasks with 300 expert demonstrations. To optimize policy with human preferences, we collect 50 trajectories per task under different seeds for RoboMimic tasks. For evaluation, we conduct 50 trials under three unseen seeds for each task, and report the average success rate.
@@ 4.2 Comparison Experiments
We compare our adaptive reweighting preference optimization method with other approaches to evaluate the effectiveness for VLA model fine-tuning. To ensure fairness, we fine-tune OpenVLA [18] for manipulation tasks as a base model and improve the base model with other comparison methods.
@@
As shown in Table 1, we first compare the behavior cloning objective methods. The results reveal that after fine-tuning with interaction data, these methods fail to outperform the base model, which demonstrates that existing behavior cloning approaches struggle to achieve efficient adaptation in large-scale VLA models. A key challenge stems from the distribution shift between expert trajectories and interaction trajectories. Without mechanisms to retain the base model’s knowledge under the standard behavior cloning objectives, this shift makes it particularly difficult for large-scale VLA models to fit the multimodal distribution of expert and interaction datasets.
@@
Among all compared preference learning based methods, DPO yields the weakest performance. This result stems from its exclusive reliance on synthetic paired failure actions for optimization, which lacks exposure to real-world errors essential for teaching robots mistake avoidance through interaction. On the other side, TPO fails to deliver stable performance gains on multiple tasks while our attains stable performance gains relative to the base model. The TPO method employs negative samples to regularize model preference alignment optimization, but introduces instability through random sampling. In contrast, our method utilizes KL divergence to estimate the mean margin between the updated model and the reference model, which not only enables more stable learning but also better preserves prior knowledge. Compared with KTO, our method leverages the adaptive reweighting method to achieve more precise control over the importance weights of both positive and negative samples, delivering more notable performance improvements.
@@
Ours 34% 62% 40% 45.3%

### v3

# Human-assisted Robotic Policy Refinement via Action Preference Optimization

@@ Abstract
Establishing a reliable and iteratively refined robotic system is essential for deploying real-world applications. While Vision-Language-Action (VLA) models are widely recognized as the foundation model for such robotic deployment, their reliance on offline expert demonstrations critically limits their capacity for post-deployment refinement. To mitigate this limitation, we introduce Action Preference Optimization (APO), a method designed to refine VLA models by human-assisted preference alignment gathered through interaction with environments. This method begins with a human-robot collaboration framework for reliable failure correction and interaction trajectory collection through human intervention. However, directly leveraging these interaction trajectories for preference optimization is non-trivial due to the challenges of irreversible robotic actions and token distribution mismatch. To solve this, APO proposes an adaptive reweighting algorithm with binary desirability signals derived from interaction, empowering VLA models effectively suppress failure-prone actions while enhancing corrective action adaptation. Ultimately, APO equips VLA models with the crucial capability to learn from failure, paving the way for their iterative refinement and reliable deployment in dynamic environments. The experiments conducted in simulation and real-world scenarios prove superior generalization and robustness of our human-assisted framework across a variety of manipulation tasks. We believe this work could bring insights for efficient and stable optimization of VLA models through human-robot collaboration.
@@ 1 Introduction
Fostering continuous improvement is crucial for the development of robust robotic manipulation systems in real-world scenarios [10, 28, 45]. Benefiting from the capacity for generalizable reasoning and scalable learning, Vision-Language-Action (VLA) models [3, 4, 8, 18, 43, 21, 47] have been widely recognized as the foundation model for such robotic deployment systems. However, prevailing training paradigm for these models hinges on large-scale datasets of expert demonstrations. This severely limits their post-deployment refinement, as they lack intrinsic ability to continually learn from failures or adapt to novel scenarios encountered in real world.
@@
To enhance the continuous learning ability of robotic systems, interactive imitation learning frameworks [7, 17] have been developed to refine error-prone trajectories via iterative human-in-the-loop correction feedback. Among them, behavior cloning [15, 22, 35] has been widely utilized to fine-tune a base policy model with manually corrected intervention data in a supervised learning fashion. In contrast, recent methods [20, 26] seek to propose effective off-policy reinforcement learning algorithms from sub-optimal human intervention trajectories. However, behavior cloning fails to fully exploit failure trajectories, which are valuable signals for learning robust policies. At the same time, reinforcement learning methods encounter significant scalability limitations in training large-scale VLA models, due to the inherent instability and challenge of developing generalizable value functions. To date, the effective adaptation of VLA models for downstream manipulation tasks remains understudied, particularly within sub-optimal human intervention paradigms.
To bridge this gap, we propose Action Preference Optimization (APO), a new paradigm moves beyond the limitations of both behavior cloning and reinforcement learning. By learning from action-level preferences captured during interactions, our approach fully exploits the valuable information in failure trajectories while maintaining the optimization stability required for large-scale VLA models.
Our method is founded on a human-robot collaboration framework designed to ensure reliable deployment while simultaneously generating data for policy refinement, as illustrated in Figure 1(a). When the robot encounters challenging situations, real-time human interventions not only guarantee successful task completion but also provide corrective trajectories. These trajectories are collected as preference pairs to refine the policy. Furthermore, to address the imbalanced distribution of action types in the collected data, we employ a balanced sampling method. This ensures a proportional representation of all interaction data for the subsequent VLA preference optimization
@@
Figure 1: Our method consists of two key components: (a) the human-robot collaboration deployment framework for reliable deployment and interaction trajectory collection with human intervention. (b) the action preference optimization process with adaptive reweighting for VLA models learning from sub-optimal interaction trajectories. The size of each circle represents its weight during training.
@@
However, directly applying this preference data to fine-tune autoregressive VLA models presents two significant, interrelated challenges: (1) Irreversible interaction: While LLMs often require paired preference data, the irreversible nature of physical interaction makes it extremely difficult to gather perfect positive-negative action samples under identical conditions. (2) Token probability mismatch: Autoregressive VLA models discretize continuous actions into tokens, causing a fundamental mismatch between token probabilities and the true action loss, which complicates preference alignment. To address these pressing problems, we first employ Kahneman & Tversky’s prospect theory [11, 40] to formulate a novel preference alignment objective that learns from binary desirability signals derived from interaction. This objective relaxes the stringent demands of preference pairs, making it suitable for learning from irreversible robotic interaction trajectories. Furthermore, we propose an adaptive reweighting method that leverages decoded continuous actions to effectively guide preference optimization in the discrete action token space. This approach addresses the challenge of action token probability mismatch via the dynamic modulation of sample-wise training weighting, thereby concentrating gradient optimization on failure-prone interaction actions. Through this targeted weight refinement, we successfully apply preference alignment optimization to VLA models, enhancing performance when learning from sub-optimal manipulation correction trajectories.
@@
To systematically evaluate the effectiveness of our proposed system, we conduct a comprehensive set of experiments in RoboMimic [29] simulation environments. The empirical results demonstrate that APO facilitates rapid adaptation in in-distribution scenarios while maintaining robust performance across a variety of unseen perturbations. Furthermore, lifelong learning experiments demonstrate the framework’s capacity for iterative improvement through human intervention. To evaluate the practical viability of the proposed framework, we conducted real-world experiments on fine-grained insertion tasks under a range of disruption conditions, demonstrating its robustness and applicability in real-world robotic manipulation scenarios.
2 Related Works
@@ 2.1 Vision-Language-Action Models
Achieving generalizable robotic manipulation remains a significant challenge within the field of robotics. Motivated by recent advances in foundation models [19, 39, 42, 32, 44, 48], some works [6, 12, 41] attempt to construct large-scale real-world robotic datasets to facilitate development of generalizable Vision-Language Action (VLA) models. Building upon these datasets, recent research [5, 18, 33] formulates robotic action prediction as a next token prediction problem within the framework of VLMs. In contrast, alternative studies [24, 38] investigate the applicability of diffusion-based methods to model multi-modal action distributions, thereby facilitating robustness in manipulation tasks. While these works focus on behavior cloning from expert demonstrations, Grape [49] proposes a trajectory-level preference alignment method to boost generalizability by incorporating both successful and failed trials. However, the requirements of paired trajectories under the same conditions make it infeasible in real-world scenarios. In this work, we propose the action preference optimization method to continuously refine VLA models by integrating human-in-the-loop intervention preference data.
@@ 2.2 Preference Alignment of Large Language Models
Contemporary methods [9, 30, 31, 50] implement Reinforcement Learning from Human Feedback (RLHF) through a two-stage method, which first trains a reward estimation model and optimizes LLMs to maximize the given estimated reward with a reinforcement learning method [13, 36, 25]. However, this paradigm is slow and unstable in practice. DPO [34] proposes a single-step alternative that reparameterizes the RLHF objective into a closed-form loss function to directly maximize the log-likelihood margin between preferred and dispreferred outputs. Extending this framework, KTO [11] introduces the human-aware losses for learning from a binary signal of whether an output is desirable, which bypasses the need for intricate preference annotation altogether. In this work, we adapt the preference alignment optimization method for Vision-Language-Action models. Through an adaptive reweighting approach, we mitigate the irreversible interactions and token probability mismatch challenges when transferring preference learning methods from LLMs to VLA models.
@@ 2.3 Human-robot Interactive Learning
Interactive imitation learning [1, 7] has been proposed to refine robot actions through human feedback. While prior research [17, 22, 35] necessitates constant human supervision to intervene in the robot’s actions, more recent studies [14, 23, 46] have introduced dynamic models for automatic failure detection and real-time monitoring. In contrast, RLIF [26] leverages human intervention signals as rewards for off-policy RL, while HIL-SERL [27] presents a human-in-the-loop, vision-based RL system tailored for dexterous manipulation tasks. However, these RL approaches encounter difficulties in large-scale VLA model training, primarily due to unstable gradient optimization. In this work, we propose the action preference optimization method to ensure the stable optimization of policies from action-level preferences captured during interaction.
3 Method
In this section, we introduce Action Preference Optimization (APO), a method designed to facilitate continuous iterative improvement of Vision-Language-Action (VLA) models. As detailed in Algorithm 1, APO aligns the model with human preferences gathered through human-robot collaboration deployment within environment.
@@ 3.1 Human-robot Collaboration Deployment
To ensure reliable deployment and interaction trajectory collection, our method is founded on the human-robot collaboration framework for real-time intervention and interaction data acquisition.
@@
Algorithm 1 Action Preference Optimization
@@
[…]
@@
[…]
@@
27:function Optimization($\pi_{\theta}$$\pi_{ref}$$\mathcal{D}_{h}$) 28:  for n gradient steps do 29:   Balanced Sample $(o_{i},a_{i},c_{i})_{i=1}^{i=B}\sim\mathcal{D}_{h}$ 30:   $l_{i}\xleftarrow{}|\pi_{\theta}(o_{i})-a_{i}|_{1}$ 31:   $w_{i}\xleftarrow{}\frac{l_{i}}{\sum_{i=1}^{i=B}l_{i}}$ 32:   if $c_{i}\neq 0$ then $\lambda_{D_{i}}=1-e^{-\beta_{D}*w_{i}}$ 33:   else $\lambda_{U_{i}}=e^{-\beta_{U}*w_{i}}$ 34:   end if 35:   $\theta^{*}=\arg\max_{\theta}\mathbb{E}[-v(o,a,\pi_{\theta},\pi_{ref},\lambda_{D},\lambda_{U})]$ 36:  end for 37:  return $\pi_{\theta^{*}}$ 38:end function
3.2 Action Preference Optimization
@@
In conclusion, we propose the action preference optimization method as demonstrated in algorithm 1. This approach leverages the human-robot collaboration deployment for reliable task execution and interaction trajectories collection, while the action preference optimization process provides stable autoregressive VLA optimization with adaptive reweighting. Through iterative human-robot collaboration deployment and action preference optimization, we could achieve continual improvement from interaction with environments for autoregressive VLA models.
@@ 4 Experiments
To comprehensively evaluate our action preference optimization method for effective downstream adaptation, we propose experiments to validate the following questions:
  • How effective is APO at promoting adaptation to in-distribution scenarios? Section 4.2
  • Does APO maintain effective learning performance in novel scenarios despite various disruptions? Section 4.3
  • Does APO demonstrate the ability to achieve iterative improvement during deployment? Section 4.4
  • How well does APO generalize to different VLA models? Scetion 4.5
  • Could APO be applied in fine-grained real-world scenarios? Section 4.6
  • To what extent does the action-level preference optimization method enable APO to learn action correction? Section 4.7
@@ 4.1 Experiment Settings
Implementation Details. In this work, we fine-tune the OpenVLA [41] model for target manipulation tasks as the base model. We employ LoRA [16] for parameter-efficient tuning, configuring rank $r=32$ with a batch size of 16 across 8 NVIDIA A100 GPUs. Further, we deploy the base model to interact with environments, where human operators perform real-time corrective interventions via a SpaceMouse device to rectify failures during execution. We set $K=10$ to identify and annotate undesirable behaviors automatically. The human-assisted interaction trajectory is shown in Figure 7, which is segmented into robotic automatic execution, the failure action, and the human intervention types by the timing of human correction. Based on the interaction trajectories collected during task execution, we fine-tune the base model $\pi_{ref}$ with our action preference optimization method, using a learning rate of $5e{\text{-}}5$ and a batch size of 8 across 4 NVIDIA A100 GPUs. To ensure the stability of preference alignment training, we employ balanced sampling to ensure that each batch contains 50% expert actions, 25% human intervention actions, and 25% failure actions.
Figure 2: The demonstration of our human-assisted interaction trajectory.
Simulation Environments Details. For a comprehensive evaluation, we validate these methods on fine-grained manipulation tasks within the RoboMimic [29] simulation environment, such as ’make coffee’ and ’toy assembly’. In the RoboMimic environment, we fine-tune the pretrained OpenVLA model for these 4 long-horizon manipulation tasks with 300 expert demonstrations. To optimize policy with human preferences, we collect 50 trajectories per task under different seeds for RoboMimic tasks. For evaluation, we conduct 50 trials under three unseen seeds for each task, and report the average success rate.
@@ 4.2 Comparison Experiments
We compare APO with other approaches to evaluate the effectiveness for VLA model fine-tuning. To ensure fairness, we fine-tune OpenVLA [18] for manipulation tasks as a base model and improve the base model with other comparison methods.
@@
As shown in Table 1, we first compare the behavior cloning objective methods. The results reveal that after fine-tuning with interaction data, these methods fail to outperform the base model, which demonstrates that existing behavior cloning approaches struggle to achieve efficient adaptation in the context of large-scale VLA models. A key challenge stems from the distribution shift between expert trajectories and interaction trajectories. Without mechanisms to retain the base model’s knowledge under the under standard behavior cloning objectives, this shift makes it particularly difficult for large-scale VLA models to effectively fit the complex, multimodal distribution arising from the combined expert and interaction datasets.
@@
Among all compared preference learning based methods, DPO yields the weakest performance. This result stems from its exclusive reliance on synthetic paired failure actions for optimization, which lacks exposure to real-world errors essential for teaching robots mistake avoidance through interaction. On the other side, TPO fails to deliver stable performance gains on multiple tasks while APO attains stable performance gains relative to the base model. The TPO method employs negative samples to regularize model preference alignment optimization, but introduces instability through random sampling. In contrast, APO utilizes KL divergence to estimate the mean margin between the updated model and the reference model, which not only enables more stable learning but also better preserves prior knowledge. Compared with KTO, APO leverages the adaptive reweighting method to achieve more precise control over the importance weights of both positive and negative samples, delivering more notable performance improvements.
@@
APO 34% 62% 40% 45.3%
Figure 4: Lifelong learning results of APO method.